Just about the time I thought I got it, here's a new one that I can't quite figure out how to proceed with.
Find f^-1(x):
f(x) = x^2 + 9 for x>= 0.
Thanks in advance for any help on this!

Answer by uma(370)

given f(x) = x^2 + 9
Let y = x^2 + 9
To find f^-1 we first interchange x and y
==> x = y^2 + 9
==> x - 9 = y^2
==> sqrt(x - 9) = y
Thus f^-1 (x) = sqrt(x - 9)